Parth (parthkohli):

User P gets 3 testimonials per week. User L gets 1 testimonial per week. User P has 45 testimonials. User L has 53. When will both of them be equal?

OpenStudy (lgbasallote):

lol bogus!!

Parth (parthkohli):

\( \color{Black}{\Rightarrow 45 + 3x = 53 + x }\) \( \color{Black}{\Rightarrow 3x = 8 + x}\) \( \color{Black}{\Rightarrow 2x = 8 }\) I don't know how to isolate.

Parth (parthkohli):

CAN YOU GUYS TELL THE ANSWER!?!?! I WILL FAIL MY ONLINE TESTS IF YOU DONT

OpenStudy (lgbasallote):

is it 3 am there and you need this homework to graduate?

Parth (parthkohli):

YES DO IT DUDE TELL TEH ANSWER

OpenStudy (lgbasallote):

and are you dyslexic and have ADHD?

Parth (parthkohli):

NO BUT I WANT YOU TO TELL THE ANSWER AND I WILL CALL THE POLICE IF U DONT

Parth (parthkohli):

@annas I am reporting you for sniping..jk

OpenStudy (lgbasallote):

oh please dont ill tell you the answer!!

OpenStudy (anonymous):

ok lol

OpenStudy (anonymous):

Is 28 is the answer???

Parth (parthkohli):

How to solve 2x = 8!?!??!?!?!

Parth (parthkohli):

I get x = 8 - 2 x = 6

OpenStudy (lgbasallote):

you divide by 2..that would give you around 6

OpenStudy (lgbasallote):

you're right!!! ^_^

OpenStudy (lgbasallote):

congrats @ParthKohli

Parth (parthkohli):

YAY 6 is correct :D

Parth (parthkohli):

Thanks lgbasallote

OpenStudy (anonymous):

28 days = 4 weeks.... So, in term of weeks 4 is the anwer...

OpenStudy (lgbasallote):

any time parth ;)

OpenStudy (anonymous):

lol wow